# Elastic Constants and Bending Rigidities from Long-Wavelength Perturbation Expansions

### v1

@@ III.1 Equations of motion for bending waves
and if the elastic stiffness tensor $C_{\alpha\beta,\gamma\delta}$ is homogeneous over the plate thickness $h$, the above integral is then reduced to $D_{\alpha\beta,\gamma\delta}=h^{3}C_{\alpha\beta,\gamma\delta}/12$. As discussed in the literature [71, 72], the validity of using such continuum approximation to calculate bending rigidities can be questionable for nanoscale materials, due to the ambiguous definition of the effective bending thickness which can be even direction-dependent in elastic anisotropic materials. Therefore, we develop in Sec. III.2 a microscopic theory of long-wavelength bending vibrations to predict the flexural properties of 2D materials.
@@
Besides, it has been also shown that there can be a strong electromechanical coupling between the strain gradient (i.e. the bending curvature) and macroscopic electric fields due to flexoelectric effects in piezoelectric 2D nanoplates [110, 111, 112], which may further alter their mechanical properties and depends on the electrical boundary conditions. The first-principles calculations of flexoelectric coefficients in 2D materials have only recently appeared [113, 114] such that we restrict ourselves to the case of short-circuit boundary conditions which corresponds to the removal of macroscopic long-range electrostatic interactions in the 2D formalism (see Sec. III.4 for details). The effect of flexoelectricity on bending rigidities is left for the future investigation.
@@ III.2 Long-wavelength flexural acoustic vibrations
As demonstrated in App. B, the solvability condition given by Eq. (123) for the fourth-order equation gives rise to the equations of motion for acoustic bending modes in LD materials as
@@
$W^{\rm{CI}}_{\alpha\beta,\gamma\delta,\lambda\mu}$ $=\sum_{\kappa}W^{\mathrm{CI},\kappa}_{\alpha\beta,\gamma\delta,\lambda\mu}=\sum_{\kappa,\kappa^{\prime}}\Phi^{(4),\gamma\delta\lambda\mu}_{\kappa\alpha,\kappa^{\prime}\beta},$ (73)
@@
$[\omega^{(2)}_{\rm ZA}(\mathbf{q})]^{2}=q_{\alpha}q_{\beta}q_{\gamma}q_{\delta}\frac{W_{zz,\alpha\beta,\gamma\delta}}{M},$ (75)
@@ III.3 Closed-form expression for bending rigidity tensor
The microscopic expression for the bending rigidity tensor of 2D materials can be obtained by connecting the equations of motion for bending waves derived from the lattice-dynamical theory and the Kirchhoff plate theory, i.e. Eqs. (71) and (75). By introducing the 2D mass density $\rho_{2D}=M/A$ with the primitive unit cell area $A$, we can obtain
$[\omega^{(2)}_{\rm ZA}(\mathbf{q})]^{2}=q_{\alpha}q_{\beta}q_{\gamma}q_{\delta}\frac{1}{\rho_{2D}}\frac{W_{zz,\alpha\beta,\gamma\delta}}{A}.$ (76)
@@
where $z$ is assumed to be the out-of-plane direction. The different definitions of bending rigidity tensor again arise the dynamic reciprocal-space versus static real-space formulations, similar to the case of strain-gradient elasticity tensor [94, 95]. The dynamic bending rigidity tensor itself is fully symmetric with respect to the exchange of any of its four indices, $\alpha$, $\beta$, $\gamma$, and $\delta$; in contrast, the static tensor introduced in Eq. (60) exhibits lower symmetry and must adhere to the same symmetry relations as the elastic stiffness tensor described in Eq. (3). While the clamped-ion part of bending rigidities $W^{\rm{CI}}_{zz,\alpha\beta,\gamma\delta}$ is inherently fully symmetric through its definition as a fourth-order derivative of the zone-center dynamical matrix, the lattice-mediated part $W^{\rm{LM}}_{zz,\alpha\beta,\gamma\delta}$ is only symmetric by exchanging indices within the two pairs $\alpha\beta$ and $\gamma\delta$, where we have already imposed the corresponding symmetrizations (see App. B). Since only symmetric parts contribute to the strain-energy density, from Eqs. (3) and (60), the symmetrized real-space bending rigidity tensor can be further introduced as
@@
which satisfies the symmetry relation of elastic stiffness tensor. The bending rigidity tensor defined in Eq. (78) is general and can also be applied for 1D materials, where the unit cell area $A$ is replaced by the unit cell length $L$ along the 1D axis.
We note again that the bending ZA modes in 2D materials can only emerge when the rotational invariance and vanishing stress conditions (also known as the equilibrium conditions) introduced in App. A are both satisfied [90, 109]. This statement agrees with a recent study [115] which shows that the rotational invariance protects the quadratic ZA phonons in the long-wavelength limit against the phonon-phonon interactions, prohibiting the divergence of bending rigidity as well as allowing the sound propagation in 2D system. Besides, our numerical simulations reveal non-zero components of the elastic tensor involving the out-of-plane direction in 2D materials when the aforementioned invariance conditions are not met, further highlighting their importance.
@@ III.4 Separation of 2D long-range electrostatic interactions
where $f(\mathbf{q})=1-\tanh(|\mathbf{q}|L/2)$ is the range separation function with $L$ denoting the characteristic length scale of the range separation, and $\bm{\mathcal{Z}}$ are the generalized effective charge tensor that can be multipolarly expanded as:
@@
where $t$ is the vacuum thickness and $\mathbb{I}$ is an identity tensor. One can find that $\bm{\epsilon}^{\infty}$ is not a well-defined quantities in 2D materials, since it depends on the vacuum thickness $t$, while $\bm{\alpha}$ is thickness-independent. Also, the definition of $\alpha_{zz}$ here differs from the one $\alpha_{zz}=\frac{t}{4\pi}(1-\epsilon^{\infty,-1}_{zz})$ that can be found in the literature [97, 123], which is due to the fact that we impose the open-circuit boundary condition in the out-of-plane direction through the 2D Coulomb truncation technique [122], and this is equivalent to multiplying $\alpha_{zz}$ by a factor of $\epsilon^{\infty}_{zz}$. As a result, one can readily use the short-circuit multipole tensors in Eq. (79) that are the standard outputs of most DFT codes. For the practical calculation of the long-range part of dynamical matrix in 2D crystals, one can still apply the Ewald summation technique described in Sec. II.4, and the range-separation function $f(\mathbf{q})\simeq 1-|\mathbf{q}|L/2$ varies linearly for small momenta, unlike a quadratic behavior of the 3D counterpart. We find that such difference in the long-wavelength limit has a significant effect on choosing a suitable range-separation parameter $L$ in 2D materials for the accurate calculation of elastic tensors. In contrast to bulk solids, the short-circuit elastic constants of 2D materials become very sensitive to $L$. In order to address this obstacle, we utilize the same strategy as proposed for the bulk case to determine the optimal $L$, which is achieved through a minimization of real-space IFCs in Eq. (56) with only the macroscopic $\mathbf{G=0}$ component removed. Rigorously, $\mathbf{G=0}$ is the only long-range component that induces the macroscopic electric field effects, and any $\mathbf{G\neq 0}$ corresponds to the short-range local field effects [124]. Lastly, it should be emphasized again that the implementation of Ewald summation with only $\mathbf{G=0}$ would lead to the real-space IFCs having some imaginary parts as a signature of the breakdown of the Fourier transform of the resulting short-range dynamical matrix. Such an issue is understood in the sense that the non-analytic correction to dynamical matrix is not only limited to the vicinity of the Brillouin zone center if only $\mathbf{G=0}$ is included in the Ewald summation; the non-analytic correction on the zone boundary also becomes unsymmetric with an even $\Gamma$-centered $\mathbf{q}$ grid, which consequently results in the presence of imaginary parts in the real-space IFCs. Therefore, the calculation of short-circuit elastic constants in 2D materials are performed with the real-space IFCs where the long-range effects are removed from a regular Ewald summation over a $\mathbf{G}$-grid, but with the range-separation parameter $L$ obtained with only $\mathbf{G=0}$ removed.
Figure 1: (a) Phonon dispersion and (b) independent components of the elastic tensor for silicon as a function of supercell size. The inset of panel (a) shows the zoom-in acoustic branches around the $\Gamma$ point along the K and L directions. The dotted lines in the panel (b) is the reference values of elastic constants from the finite-difference calculations using the thermo_pw code [55].
@@ IV Results and discussion
To validate our new method for computing elastic constants in Sec. II, we benchmark eight representative solids which have been widely explored by both theoretical calculations and experimental measurements: silicon, NaCl, GaAs and rhombohedral BaTiO3 are chosen as the bulk materials, while monolayer graphene, h-BN, MoS2 and InSe are selected for 2D materials. We also compute and validate the bending rigidities of graphene, h-BN, MoS2 and InSe. Specifically, in the case of bulk system, silicon is chosen as an infrared-inactive material; NaCl is selected as an infrared-active solid without quadrupoles, while GaAs is infrared-active with also the contribution from quadrupoles. Finally, BaTiO3 is chosen as a representative example of a low-symmetry system. Moreover, graphene is chosen as a prototype of 2D system, and h-BN is an example for the infrared-active 2D materials; MoS2 represents the well-known 2D family of transition metal dichalcogenides, while InSe is selected for another 2D class of III–VI metal chalcogenides. The calculation details of our first-principles simulations are summarized in App. F, except for h-BN, MoS2 and InSe where we have used the computational parameters from Ref. [98].
Table 1: The dynamical Born effective charge $Z$ [$e$], quadrupole $Q$ [$e\cdot\mathrm{bohr}$], octupole $O$ [$e\cdot\mathrm{bohr}^{2}$], clamped-ion dielectric permittivity $\bm{\epsilon}^{\infty}$, and dielectric dispersion tensors $\bm{\epsilon}^{(4)}$ [bohr2] tensors of Si, NaCl and GaAs. The values in the column this work are obtained via a least-square fit to the Taylor expansion of unscreened charge density response in Eq. (43) and dielectric screening function in Eq. (II.4), compared with those calculated from DFPT. Only symmetry-independent components are shown. The DFPT quadrupole values for silicon and GaAs are taken from Ref. [125].
@@
$\epsilon^{(4)}_{xxxx}$172.780
$\epsilon^{(4)}_{xxyy}$90.933
@@
$\epsilon^{(4)}_{xxxx}$2.353
$\epsilon^{(4)}_{xxyy}$1.184
@@
$\epsilon^{(4)}_{xxxx}$241.495
$\epsilon^{(4)}_{xxyy}$202.947
Table 2: Theoretical and experimental elastic constants, bulk ($K$), shear ($G$) and Young’s ($E$) moduli [GPa], Poisson’s ratio ($\nu$) and longitudinal ($v_{\rm l}$) and transverse ($v_{\rm t}$) sound velocities [m/s] of silicon, NaCl and GaAs. For this work, a $14\times 14\times 14$ supercell is adopted to obtain the elastic tensor of silicon and GaAs, while a $12\times 12\times 12$ supercell is used for NaCl. The results of thermo_pw are from the finite-difference of strain-stress relation. The superscript “CI” denotes the clamped-ion approximation, and the relaxation of internal atomic positions is only allowed in silicon and GaAs for their $C_{44}$ component. To obtain the elastic moduli, the polycrystalline average based on the Voigt-Reuss-Hill scheme is utilized. The experimentally measured Young’s modulus, Poisson’s ratio and speed of sound for GaAs are single-crystal results, along the [100] crystallographic direction.
@@
Figure 2: (a) Phonon dispersion with zoom-in inset and (b) independent components of elastic tensor for NaCl as a function of supercell size where the standard (DD) long-range has been removed. In the panel (b), the elastic constants of NaCl obtained without the long-range removal are shown in the inset, and the dotted lines are the reference values from the thermo_pw finite-difference calculations.

### v2

@@ III.1 Equations of motion for bending waves
and if the elastic stiffness tensor $C_{\alpha\beta,\gamma\delta}$ is homogeneous over the plate thickness $h$, the above integral is then reduced to $D_{\alpha\beta,\gamma\delta}=h^{3}C_{\alpha\beta,\gamma\delta}/12$. As discussed in the literature [77, 78], the validity of using such continuum approximation to calculate bending rigidities can be questionable for nanoscale materials, due to the ambiguous definition of the effective bending thickness which can be even direction-dependent in elastic anisotropic materials. Therefore, we develop in Sec. III.2 a microscopic theory of long-wavelength bending vibrations to predict the flexural properties of 2D materials. We further note that the bending rigidity tensor defined in this work is the same as the one entering the expansion of the energy density in terms of curvatures, as explained in Eq. (S41) of the SM [118], allowing for the connection between microscopic and macroscopic elasticity theory (see Chapter 7 of Ref. [117] for further details and derivations).
@@
Besides, it has been also shown that there can be a strong electromechanical coupling between the strain gradient (i.e. bending curvature) and macroscopic electric fields due to flexoelectric effects in piezoelectric 2D nanoplates [120, 121, 122], which may further alter their mechanical properties and depends on the electrical boundary conditions. The first-principles calculations of flexoelectric coefficients in 2D materials have only recently appeared [123, 124] such that we restrict ourselves to the case of short-circuit boundary conditions, which corresponds to the removal of macroscopic long-range electrostatic interactions in the 2D formalism (see Sec. III.4 for details). The effect of flexoelectricity on bending rigidities is left for future investigation.
@@ III.2 Long-wavelength flexural acoustic vibrations
As demonstrated in Appendix B, the solvability condition given by Eq. (123) for the fourth-order equation gives rise to the equations of motion for acoustic bending modes in LD materials as
@@
$W^{\rm{CI}}_{\alpha\beta,\gamma\delta,\lambda\mu}$ $=\sum_{\kappa}W^{\mathrm{CI},\kappa}_{\alpha\beta,\gamma\delta,\lambda\mu}=\frac{1}{24}\sum_{\kappa,\kappa^{\prime}}\Phi^{(4),\gamma\delta\lambda\mu}_{\kappa\alpha,\kappa^{\prime}\beta},$ (73)
@@
$\omega^{(2)}_{\rm ZA}(\mathbf{q})]^{2}=q_{\alpha}q_{\beta}q_{\gamma}q_{\delta}\frac{W_{zz,\alpha\beta,\gamma\delta}}{M},$ (75)
@@ III.3 Closed-form expression for bending rigidity tensor
The microscopic expression for the bending rigidity tensor of 2D materials can be obtained by connecting the equations of motion for bending waves derived from the Kirchhoff plate theory and the lattice-dynamical theory, i.e. Eqs. (63) and (75). By introducing the 2D mass density $\rho_{2D}=M/A$ with the primitive unit cell area $A$, we can obtain
$\omega^{(2)}_{\rm ZA}(\mathbf{q})]^{2}=q_{\alpha}q_{\beta}q_{\gamma}q_{\delta}\frac{1}{\rho_{2D}}\frac{W_{zz,\alpha\beta,\gamma\delta}}{A}.$ (76)
@@
where $z$ is assumed to be the out-of-plane direction. The different formulations of bending rigidity tensor again stem from its definition in reciprocal or real space, similar to the case of strain-gradient elasticity tensor [101, 102]. In general, the dynamic bending rigidity tensor from the reciprocal-space formulation is totally symmetric with respect to any permutation of its four indices, $\alpha$, $\beta$, $\lambda$, and $\delta$. However, this is only true within the clamped-ion approximation, since the pure electronic contribution $W^{\rm{CI}}_{zz,\alpha\beta,\gamma\delta}$ is inherently fully symmetric through its definition as a fourth-order derivative of the zone-center dynamical matrix. When the lattice-mediated part $W^{\rm{LM}}_{zz,\alpha\beta,\gamma\delta}$ is included, the resulting total tensor is only symmetric upon exchanging indices within the two pairs $\alpha\beta$ and $\gamma\delta$, if $W^{\rm{LM}}_{zz,\alpha\beta,\gamma\delta}$ is properly symmetrized [see Eqs. (119) to (122) and the related discussion in Appendix B]. This can be further understood by noticing that $W^{\rm{LM}}_{zz,\alpha\beta,\gamma\delta}$ is a mixed derivative indicated by the internal strain effects. On the other hand, using the real-space formulation, the static bending rigidity tensor introduced in Eq. (60) must adhere to the same symmetry relations as the elastic stiffness tensor described in Eq. (3). Thus, by imposing the additional symmetry $\alpha\beta\leftrightarrow\gamma\delta$, the symmetrized bending rigidity tensor can be further introduced as
@@
which satisfies the symmetry relation of elastic stiffness tensor and is consistent with the fact that only the symmetric parts contribute to strain-energy density. The applicability of a dynamic theory to study the static response is based on the establishment of mechanical equilibrium conditions between the applied external load and the internal strain response. Although individual atomic components of the bending rigidity tensor are dynamic, their overall contribution becomes static, similar to the case of strain-gradient elasticity [101]. The bending rigidity tensor defined in Eq. (78) is expected to be general and should also be applicable for one-dimensional (1D) materials, where the unit cell area $A$ is replaced by the unit cell length $L$ along the 1D axis.
Importantly, as a perturbative approach, our obtained close-form expression for the fourth-rank bending rigidity tensor in Eq. (78) adopts only the primitive unit cell of 2D materials, which avoids adding an explicit bending curvature to monolayer structures and allows the accurate calculations of bending rigidities in the perturbative limit. These advantages make our approach more favorable to the supercell method [82] and cyclic DFT calculations [83, 84, 89], which both fit the total energy of bent systems to curvature radii, requiring careful tests of radius range and magnitude to ensure the accuracy and convergence of bending rigidity in the small curvature limit. Although an implementation of DFPT into cyclic DFT would also allow for the calculation of bending rigidities, the recent DFPT development of cyclic DFT is focused only on phonon frequencies due to its complexity [125]. In addition, we note our bending rigidity formalism is only exact in the linear asymptotic limit where non-linear effects are neglected, and the bending ZA modes in 2D materials can only emerge when the rotational invariance and vanishing stress conditions (also known as the equilibrium conditions) introduced in Appendix A are both satisfied [97, 119]. This statement agrees with a recent study [126] which shows that the rotational invariance protects the quadratic ZA phonons in the long-wavelength limit against the phonon-phonon interactions, prohibiting the divergence of bending rigidity as well as allowing the sound propagation in 2D system. Besides, our numerical simulations reveal non-zero components of the elastic tensor involving the out-of-plane direction in 2D materials when the aforementioned invariance conditions are not met, further highlighting their importance.
@@ III.4 Separation of 2D long-range electrostatic interactions
where $f(\mathbf{q})=1-\tanh(|\mathbf{q}|L/2)$ is the range separation function with $L$ denoting the characteristic length scale of the range separation, and $\bm{\mathcal{Z}}$ are the generalized effective charge tensor that can be multipolarly expanded as
@@
where $t$ is the vacuum thickness and $\mathbb{I}$ is an identity tensor. One can find that $\bm{\epsilon}^{\infty}$ is not a well-defined quantity in 2D materials, since it depends on the vacuum thickness $t$, while $\bm{\alpha}$ is thickness-independent. Also, the definition of $\alpha_{zz}$ here differs from the one $\alpha_{zz}=\frac{t}{4\pi}(1-\epsilon^{\infty,-1}_{zz})$ that can be found in the literature [104, 134], which is due to the fact that we impose the open-circuit boundary condition in the out-of-plane direction through the 2D Coulomb truncation technique [133], and this is equivalent to multiplying $\alpha_{zz}$ by a factor of $\epsilon^{\infty}_{zz}$. As a result, one can readily use the short-circuit multipole tensors in Eq. (79) that are the standard outputs of most DFT codes. For the practical calculation of the long-range part of dynamical matrix in 2D crystals, one can still apply the Ewald summation technique described in Sec. II.4, and the range-separation function $f(\mathbf{q})\simeq 1-|\mathbf{q}|L/2$ varies linearly for small momenta, unlike a quadratic behavior of the three-dimensional (3D) counterpart. We find that such a difference in the long-wavelength limit has a significant effect on choosing a suitable range-separation parameter $L$ in 2D materials for the accurate calculation of elastic tensors. In contrast to bulk solids, the short-circuit elastic constants of 2D materials become very sensitive to $L$. In order to address this obstacle, we utilize the same strategy as proposed for the bulk case to determine the optimal $L$, which is achieved through a minimization of real-space IFCs in Eq. (56) with only the macroscopic $\mathbf{G=0}$ component removed. Rigorously, $\mathbf{G=0}$ is the only long-range component that induces the macroscopic electric field effects, and any $\mathbf{G\neq 0}$ corresponds to the short-range local field effects [135]. Lastly, it should be emphasized again that the implementation of Ewald summation with only $\mathbf{G=0}$ would lead to the real-space IFCs having some imaginary parts as a signature of the breakdown of the Fourier transform of the resulting short-range dynamical matrix. Such an issue is understood in the sense that the non-analytic correction to dynamical matrix is not only limited to the vicinity of the Brillouin zone center if only $\mathbf{G=0}$ is included in the Ewald summation; the non-analytic correction on the zone boundary also becomes unsymmetric with an even $\Gamma$-centered $\mathbf{q}$ grid, which consequently results in the presence of imaginary parts in the real-space IFCs. Therefore, the calculation of short-circuit elastic constants in 2D materials are performed with the real-space IFCs where the long-range effects are removed from a regular Ewald summation over a $\mathbf{G}$-grid, but with the range-separation parameter $L$ obtained with only $\mathbf{G=0}$ removed.
@@ IV Results and discussion
To validate our new method for computing elastic constants in Sec. II, we benchmark eight representative solids, which have been widely explored by both theoretical calculations and experimental measurements: silicon, NaCl, GaAs, and rhombohedral BaTiO3 are chosen as the bulk materials, while monolayer graphene, h-BN, MoS2, and InSe are selected for 2D materials. We also compute and validate the bending rigidities of graphene, h-BN, MoS2, and InSe. Specifically, in the case of bulk system, silicon is chosen as an infrared-inactive material; NaCl is selected as an infrared-active solid without quadrupoles, while GaAs is infrared-active with also the contribution from quadrupoles. Finally, BaTiO3 is chosen as a representative example of a low-symmetry system. Moreover, graphene is chosen as a prototype of the 2D system, and h-BN is an example for the infrared-active 2D materials; MoS2 represents the well-known 2D family of transition metal dichalcogenides, while InSe is selected for another 2D class of III-VI metal chalcogenides. The calculation details of our first-principles simulations are summarized in Appendix F, except for h-BN, MoS2 and InSe where we have used the computational parameters from Ref. [105].
Figure 1: (a) Phonon dispersion and (b) independent components of the elastic tensor for silicon as a function of supercell size. The inset of panel (a) shows the enlarged acoustic branches around the $\Gamma$ point along the K and L directions. The dotted lines in the panel (b) are the reference values of elastic constants from the finite-difference calculations using the thermo_pw code [59].
Table 1: The dynamical Born effective charge $Z$ [$e$], quadrupole $Q$ $e~\mathrm{bohr}$], octupole $O$ $e~\mathrm{bohr}^{2}$], clamped-ion dielectric permittivity $\bm{\epsilon}^{\infty}$, and dielectric dispersion tensors $\bm{\epsilon}^{(4)}$ [bohr2] tensors of Si, NaCl and GaAs. The values in the column this work are obtained via a least-squares fit to the Taylor expansion of unscreened charge density response in Eq. (43) and dielectric screening function in Eq. (II.4), compared with those calculated from DFPT. Only symmetry-independent components are shown. The DFPT quadrupole values for silicon and GaAs are taken from Ref. [136].
@@
$\epsilon^{(4)}_{xxxx} -$172.780
$\epsilon^{(4)}_{xxyy} -$90.933
@@
$\epsilon^{(4)}_{xxxx} -$2.353
$\epsilon^{(4)}_{xxyy} -$1.184
@@
$\epsilon^{(4)}_{xxxx} -$241.495
$\epsilon^{(4)}_{xxyy} -$202.947
Table 2: Theoretical and experimental elastic constants ($C_{ij}$), bulk ($K$), shear ($G$), and Young’s ($E$) moduli [GPa], Poisson’s ratio ($\nu$), and longitudinal ($v_{\rm l}$) and transverse ($v_{\rm t}$) sound velocities [m/s] of silicon, NaCl and GaAs. For this work, a $14\times 14\times 14$ supercell is adopted to obtain the elastic tensor of silicon and GaAs, while a $12\times 12\times 12$ supercell is used for NaCl. The results of thermo_pw are from the finite-difference of strain-stress relation. The superscript “CI” denotes the clamped-ion approximation, and the relaxation of internal atomic positions is only allowed in silicon and GaAs for their $C_{44}$ component. To obtain the elastic moduli, the polycrystalline average based on the Voigt-Reuss-Hill scheme is utilized. The experiment data for Si and NaCl were measured at 298 and 300 K, respectively, while those for GaAs were extrapolated to zero temperature. The experimental Young’s modulus, Poisson’s ratio, and speed of sound for GaAs are single-crystal results, along the [100] crystallographic direction.
@@
Figure 2: (a) Phonon dispersion with enlarged inset around the $\Gamma$ point and (b) independent components of elastic tensor for NaCl as a function of supercell size where the standard ($D$) long-range has been removed. In the panel (b), the elastic constants of NaCl obtained without the long-range removal are shown in the inset, and the dotted lines are the reference values from the thermo_pw finite-difference calculations.